Fermi National Laboratory
 
Operation Summary for April 18, 2012 - April 20, 2012
 

Wednesday April 18

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), with the SeaQuest experiment off for a Neutrino Area enclosure N01 access, and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, N01 came out of access and SeaQuest resumed taking beam at 8:18 AM. Operations disabled the SeaQuest beam at 9:50 AM for accesses into Enclosures C, D, and E, G2, and NM4. Muon Ring personnel began a study at 10:22 AM. Operations received an Operating Note allowing a SeaQuest intensity readback study, in which some intensity and beam monitors were moved around; this study lasted about 45 minutes. Operations established beam to MTA at 12:28 PM. Operators had the Neutrino Area enclosures ready for access by 12:33 PM; the accesses lasted about an hour. The Muon Ring study ended at 2:31 PM. The NuMI beam permit tripped off at 2:47 PM due to a power supply’s output (E:QF107) drifting low; everything reset.

During the evening shift, at 4:07 PM, the NuMI beam permit tripped off due to QF107; everything reset. Operations received the Operational Readiness Clearance ( ORC) for the Meson FTBF experiment T-1015 at 4:11 PM. At 4:25 PM, Operations received an Operating Note that allowed SeaQuest to take low intensity beam for a study; the study began at 4:50 PM. At 6:12 PM, the NuMI beam permit tripped off when a power supply (E:HV101) tripped off; everything reset. At 6:49 PM, Operations reported a MECAR regulation problem, which held off beam for about two and a half hours. Operations disabled beam to MTA at 7:25 PM. Operations reported that T-1015 took about 20 minutes of beam to test out their equipment.

Thursday April 19

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at 2:21 AM, the NuMI beam permit tripped off due to HV101; everything reset. At 4:23 AM, Operations completed a Booster injection tune. A Muon Ring person began a study at 5:23 AM; he finished his study at 7:18 AM.

During the day shift, SeaQuest reported at 8:12 AM that they had completed their low intensity study; they resumed normal running. Operators prepared the Meson M01 enclosure for access at 9 AM. At 9:08 AM, Operations turned off all beam and loaded an energy conservation lime line into the Time Line Generator (TLG) primarily for an access into Transport for vacuum work; Linac personnel then used the time to change out a power amplifier for LRF3. Operators prepared NuMI, the Neutrino Area, and SeaQuest for access. Linac established beam to their dump at 11:42 AM. The Neutrino Area came out of access at 11:51 AM. Operations established beam to MiniBooNE at 12:15 PM. Operations established beam to the Switchyard (SY) dump at 12:20 PM. Operations established beam to MTA and SeaQuest at 12:28 PM. Operations established beam to T-1015 at 1:32 PM. A MI person began a g-2 (g minus 2) study. At 1:40 PM, the NuMI beam permit tripped off due to QF107; everything reset.

During the evening shift, 4:24 PM, the Muon Ring AP2 line ion pumps all tripped off for the first of eight times in six and a half hours; they all reset. The g-2 study ended at 5:02 PM. At 5:10 PM, BRF10 tripped off and would not reset; Operators bypassed the station and called in RF technicians to disconnect the station from the anode power supply, which took about 40 minutes. BRF11 tripped off at 5:39 PM for the first of two times in 19 minutes; it reset. The MI South Anode Power Supply tripped off at 6:47 PM on an interlock fault; it reset, but MIRF2, which also had tripped off, would not reset remotely. Operators had to reset MIRF2 locally.

Friday April 20

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest and T-1015.

During the midnight shift, at 12:33 AM, BRF11 tripped off for the first of five times in about 30 minutes; Operations reported that these trips were due to Booster transition tuning. Operations completed a 750 KeV line tune at 5:35 AM. Operations completed a 400 MeV area tune at 6:07 AM. At 7:59 AM, Operations prepared the Neutrino Area and SeaQuest for access.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo agrees with Booster, MI, and Muon Ring study plans for today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/20/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory